OpenStudy (anonymous):

which expression is not equivalent to the other expressions? 7 + 0.5 x 4^2 2.5 x (16-2^2) (3^3 + 3) ÷ 2 20 + 5^2) ÷ 3

OpenStudy (anonymous):

The first one is 7 + 0.5*16, which = 7 + 8 which = 15. The second one is 2.5*(14^2) which is over 400, way off. The third one is (27+3)/2, which = 30/2, which is 15. he last one is (20 + 25)/3, which = 45/3, which equals 15. So the answer is the second one :) Hope that helped!

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

No problem, anytime! :)